OpenStudy (anonymous):

Evaluate 27 × a-2 for a = 3. one-third one-ninth 9 3

OpenStudy (imstuck):

The negative exponent requires you to move the constant to a denominator to make it a positive exponent. \[3^{-2}=\frac{ 1 }{ 3^{2} }\]So if you multiply, you get \[27\times \frac{ 1 }{ 9 }\]Can you handle it from there or no?

OpenStudy (anonymous):

THANK YOUUUUUUUUUU

OpenStudy (imstuck):

My pleasure!